A friend of mine is having trouble with doing this problem and I have no clue how to help her out. Her teacher is telling her she to take another look at her overtime pay rate calculation..please do help her and me. thanks.
She chose option 1: Method 1:(total regular hours X hourly rate) + (overtime hours X overtime rate)=gross pay
And this is her information :
Pay rate: 14.26
Regular Hours: 80
Overtime hours: 6
Overtime rate: 1.5 (time and a half)
(80 x 14.26) + (6 x 1.5) = gross pay
80 x 14.26 = 1140.80 6 x 1.5 = 9
1140.80 + 9 = 1149.80

Found 2 solutions by Alan3354, jim_thompson5910:
Answer by Alan3354(69443)

She chose option 1: Method 1:(total regular hours X hourly rate) + (overtime hours X overtime rate)=gross pay
And this is her information :
Pay rate: 14.26
Regular Hours: 80
Overtime hours: 6
Overtime rate: 1.5 (time and a half)
(80 x 14.26) + (6 x 1.5) = gross pay *************** OT is 1.5 times $14.26, not $1.50 per hour.
---------------------
80 x 14.26 = 1140.80 6 x 1.5 = 9
1140.80 + 9 = 1149.80

Answer by jim_thompson5910(35256)

It should be

(total regular hours X hourly rate) + (overtime hours X overtime rate)=gross pay

(80 x 14.26) + (6 x 1.5 x 14.26) = gross pay

Since the overtime rate is 1.5 x 14.26 (and not just 1.5)

Notice the addition in red